MEDICATION CALCULATION #6

Order: Methyldopa (Aldomet) 0.3g, PO, bid.

Available: Methyldopa 250mg tablets.

How many tablet(s) should you give per dose?

Please show detailed calculation.
Answer #1

Order: 0.3 gram tablet

That is 300mg

Because 1gram = 1000mg

Available= 250mg tablet

1 tablet = 250mg

1/2 tablet= 125mg

1/4 tablet= 62.5mg

1/8 tablet= 31.25

1/16 tablet= 15.75

1/24 tablet = 50mg

So give 1 tablet and 1/24 tablet to reach

300mg( 250mg+50mg)